Minutes of the Tracker Technical Meeting

August 7, 2002


Action Items:

1. Luisella: review the drawings from SLAC asap.


Tray panels: Plyform is waiting for the new set of tooling.  At present there still are only 4 trays built, plus the tray with intentional defects.  An very high frequency ultrasound system in France was able to see the defects.  Now they will try the same with an interferometric system.  Plyform is closing for 3 weeks.  They will start up again August 26.  Luisella will return August 19 and will finalize the tungsten and kapton assembly systems.  Sandro thinks they will be ready the second week of September.  She will also check the top/bottom fixture drawings from SLAC, which will be time critical by that point. 


Ossie expressed some concern about the effect of water on the tray in the ultrasound tests.  The tests cannot be done in air because that gives too large an impedance mismatch.  Sandro in principle prefers the interferometric system, which can give more information and doesn’t require dipping in liquid.


Tom reported that Hytec will image damaged areas of the bottom tray with their 3-D x-ray system.  They are also cutting off a sample of carbon-carbon material to send to GSFC for analysis.


Tom spoke with Teledyne.  They verified that they can wire bond on boards from Young Electronics.  Two non-functional MCMs will be ready early to mid next week.  Robert and Tom will travel to Teledyne August 15 to review progress.


Sandro went to G&A on Monday.  He sent to us data on ladders without encapsulation (see LAT-TD-879 and LAT-TD-880).  They find excellent alignment.  Depletion voltage looks good; they sorted wafers by depletion voltage for assembly to make the ladders uniform.  There was no change after assembly.  Leakage current increased higher than in prior tests, by a factor of 1.8 on average.  This factor is the worst in the first ladders, for which the worst wafers were used.  One ladder had 10uA.  They inspected and found a dot sitting between guard and bias rings.  They removed it with alcohol and the current went down to 3uA.  G&A thinks the wafers are quite dirty.  Dust is coming from the cardboard foils in the envelopes.  The INFN groups will have to clean the wafers for G&A.  They did AC tests on 20 ladders, with a 100V step.  They found 4 additional broken caps, beyond the HPK known bad strips.  G&A now is waiting to get back clean wafers. 


Tom said that they had a successful discussion in Los Alamos of the modeling effort.  They came up with a plan to complete the analysis.  Tim Thurston will write up a report.  Tom is also working together with Martin's group on the attachment of the flexures to the Grid.